solve the quadratic equation by using the square root property.
4(a-5)^2=1

Answer by Theo(13342)

4*(a-5)^2 = 1

divide both sides by 4 to get:

(a-5)^2 = 1/4

take square root of both sides to get

a-5 = +/- sqrt(1/4)

add 5 to both sides to get:

a = 5 +/- sqrt(1/4)

since sqrt(1/4) = (1/2), equation becomes:

a = 5 +/- (1/2)

this means:

a = 5.5
or:
a = 4.5

substitute in original equation to confirm this is good.

4*(a-5)^2 = 1

if a = 4.5, this becomes 4*(-.5)^2 = 1 which becomes 4*(.25) = 1 which becomes 1 = 1 which is good.

if a = 5.5, this becomes 4*(.5)^2 = 1 which becomes 4*(.25) = 1 which becomes 1 = 1 which is also good.

answer is confirmed.

answer is:

x = 4.5
or:
x = 5.5